OpenStudy (anonymous):

p=2(l=w) solve for l

OpenStudy (anonymous):

l-w or l=w?

OpenStudy (anonymous):

oops i ment (l+w)

OpenStudy (anonymous):

Most likely minus. If that is the case, then the answer is (p+2w)/2

OpenStudy (anonymous):

(p-2w)/2 if that is the case

OpenStudy (anonymous):

\[p = 2\left( l + w \right)\] \[p = 2l + 2w\] \[p - 2w = 2l\] \[\frac{ p-2w }{ 2 } = l\]

OpenStudy (anonymous):

If you wish, you can explain in detail each step that I took within each step.

OpenStudy (anonymous):

x-y=7 solve for y

OpenStudy (anonymous):

\[x - y = 7\] \[-y = x + 7\] \[y = -\left( x+7 \right)\] \[y = -x -7\]

OpenStudy (anonymous):

thank you so much! i have more tho lol

OpenStudy (anonymous):

Apply the same principles that I showed you with the past two problems to those. I do not wish to solve everything for you, then you don't learn anything. I can give a you a few examples to get started, but you have to understand what I did in those examples, so you can do them yourself on your homework, as well as on your test. Especially your test, as I will not be there to help you there.

OpenStudy (anonymous):

i dont know hwo to do this problem. -2x+5x-9=3(x-4)-5

OpenStudy (anonymous):

I'd be glad to assist you with this one. For this, you are solving for x. I will explain each step thoroughly. -2x + 5x - 9 = 3(x-4) - 5 First of all, you distribute everything in the parenthesis. -2x + 5x - 9 = 3x - 12 - 5 From there, you will want to combine like terms on each side. 5x - 2x = 3x and -12 - 5 = -17 so 3x - 9 = 3x - 17 Now from here, get the constants(non-variable numbers on one side) 3x = 3x -17 + 9 3x = 3x - 8 Now it gets a bit interesting here. If you try to get all the x variables on one side, they seem to cancel each other out! 3x - 3x = -8 but 3x - 3x = 0, no? Therefore, this question has no answer. The answer does not exist.